A body of mass m acted on by two forces F1 and F2. If m=5.2kg, F1=3.7N and F2 =4.3N. F1 is in upward ( North ) direction and F2 is toward right ( East ) . Angle between F1 and F2 is 90° . Calculate vector acceleration of the body.

Grade:12th pass

1 Answers

Piyush Kumar Behera
7 years ago
Ok ,in this question find the net acceleration in the north direction that is F1/m=3.7/5.2=0.711similarly find the net acceleration in the east dithat is that is F2/m=4.3/5.2=0.826.now if you know vectors then as north and east direction are perpendicular so the net magnitude is√(0.711^2+0.826^2)=1.089(approx)this is from the formula of net magnitude of two vectors.Its direction is tan^-1(3.7/4.3) north of east.hope that helps !!please approve this solution to encourage me and if you don`t get the solution then please post it again here.. thanks